A machine was purchased on 1 st april, 2008. The balnce of his machine on 31 st march ,2011 is ₹ 5,83,200. Depreciation is charged @ 10% p.a. on written down value method. What was the cost price of the machine on 1 st apirl,2008 ?
1
Expert's answer
2020-10-05T16:28:58-0400

solution

"period, n = 3 \ years""residual, r = 583,200""depreciation, d= 0.01"

"Residual, r = Initial\ cost, C (1-depreciation, d)^n""583,200 = C*(0.9)^3""C =\frac{583,200}{0.9^3} = 800,000"

Answer: The cost of the machine was 800,000